OpenStudy (anonymous):

what three consecutive odd numbers have a sum of 45?

OpenStudy (anonymous):

x+x+1+x+2=45 3x+2=45 subtract 2 from each side 3x=43 divide each side by 3 x=14.3 or 14 1=14 plus 1 2=15 plus 2 3=17 < the answer

OpenStudy (richyw):

\[(x-2)+x+(x+2)=45\]\[3x=45\]\[x=15\] so your numbers would be \(13,15,17\)